Q- Suppose the government decides to increase taxes by $40 billion to increase Social Security benefits by the same amount. By how much will this combined tax transfer policy affect aggregate demand at current prices if the MPC is 0.9
Answer #1

We will have to calculate the tax multiplier and the government multiplier

Tax multiplier will be -MPC / 1-MPC = -0.9 / 0.1 = -9. and the government multiplier will be 1/ 1-MPC = 1 / 0.1 = 10.

So, the taxes will decrease the GDP by $40 x 9 = $360 and the government expenditure will increase the GDP by 400 billion.

Total aggregate demand will increase by $40 billion.